# Data Augmentation is Dead, Long Live Data Augmentation

### v1

@@ 5.   Results
We first show the inefficiency of the fine-tuning protocol in past DA papers by reporting the results they obtain without data augmentation and comparing it to our results, without data augmentation as well. We attempt to replicate the starting sizes, changing only the fine-tuning protocol to see the difference a longer patience, label smoothing, and better fine-tuning does.
@@
Kumar + Uabni TREC6 79.4 91.6
@@
Table 2 presents the results of our experiments compared to results from the literature. As we can see, having a larger patience parameter and applying label smoothing increase the performance significantly, gaining between 2 and 10%. This is the core of our argument, as our hypothesis is that the gains observed in previous studies stems only from data augmentation, giving more time for the model to learn correctly. Furthermore, if the DA strategies involve simple transformations of the sentences, it is likely that BERT is not learning new information from these transformations, and that the same result will be achieved with the Copy strategy.
@@
We now turn to DA algorithms, showing that using the updated training protocol does not result in higher gain, and that indeed the supposed gains from DA that were shown in the literature can more likely be explained by inadequate training of the classifier. Results are presented in Table 3.
@@
For the small data setting, we notice that, on average, only two strategies really outperforms the baseline : zero-shot and 3-shot generation, both using VLLMs. Of these two strategies, the 3-shot generation outperforms the zero-shot one, but not on all datasets. We explore this phenomenon in Section 6.2.
@@
An important thing to note is that while some algorithms perform statistically better than the baseline at times, so does the “Copy” strategy, which is a simple duplication of the data. As the duplication of data should bring no new information, this suggests that the training protocol was suboptimal and that the difference between the baseline and the DA results is due to this rather than a real effect of the DA algorithm. One avenue that could be explored to bridge the gap would be to use sampling with replacement while fine-tuning BERT, which should be equivalent to duplication of data [^6] There would still be minor differences as sampling with replacement would not represent each sample equally, but this difference should reasonably not have a large impact.. The results for medium sizes (500 and 1000) are presented in Appendix B, and are concordant with more recent literature showing a lack of improvement on such a setting.
@@
We give in Figure 4 the results of the statistical tests. To simplify reading, we regroup together all tests between the same algorithms, meaning all tests between algorithm A and algorithm B are compiled together. Each entry in the heatmap is therefore composed of 10 t-tests (5 datasets * 2 starting sizes), and we report the percentage of t-tests for which distributions were found to be statistically different, counting only the entries for which the row algorithm performs better than the column algorithm. This gives, for each row, the direct measure of how many times it beats the algorithm in the column by a statistically significant margin. [^7] We provide in the code the full saved results for all experiments, as well as the code necessary to run all statistical tests.
Baseline
Perfect
Copy
EDA
AEDA
BT
CBERT
CBART
T5
GPT3.5-P
GPT3.5-ZS
GPT3.5-3S
Llama2-P
Llama2-ZS
Llama2-3S
Baseline
0
20
10
10
10
30
10
10
10
20
20
40
10
30
30
Perfect
70
0
70
70
70
70
70
70
70
70
70
80
80
80
80
Copy
20
10
0
0
10
20
10
0
0
30
10
40
20
20
30
EDA
20
10
20
0
10
20
0
0
0
10
20
40
20
10
20
AEDA
10
10
0
0
0
20
0
0
0
10
0
30
20
0
0
BT
20
10
0
0
10
0
10
10
0
10
10
30
10
10
20
CBERT
20
10
10
0
10
10
0
10
0
10
0
20
10
0
20
CBART
10
10
0
0
0
10
0
0
0
10
0
30
10
10
10
T5
20
10
0
0
0
20
10
0
0
10
0
40
10
10
20
GPT3.5-P
30
10
10
20
10
40
20
20
10
0
10
30
10
20
20
GPT3.5-ZS
70
10
60
50
50
50
60
50
50
60
0
40
60
50
30
GPT3.5-3S
60
10
60
60
60
60
60
60
60
60
40
0
60
60
50
Llama2-P
20
10
10
0
10
20
20
10
0
0
0
30
0
0
10
Llama2-ZS
30
10
20
20
20
30
20
20
20
30
0
20
40
0
30
Llama2-3S
40
0
40
40
40
40
40
40
40
40
20
20
40
40
0
@@
From the statistical tests, we note that only three methods outperform the others by a significant and consistent margin : the perfect strategy, in which we simulate collecting and annotating external data, and the zero/k-shot generation with ChatGPT. All others perform equivalently and do not beat the baseline most of the time, or at least not by a significant margin. This suggests that the slight random variations we observe, which could be taken as a sign that some algorithm performs better than other, are due to randomness of the network and not a superior performance.
6.   Analysis
The preceding section evaluated data augmentation algorithms using a standard training protocol. While our main result is that the efficiency of DA is generally over-estimated, this does not imply data augmentation is never useful. In this section, we aim to test more realistic settings with the validation data, inspired by Kim et al. (2022); Zhu et al. (2023), who study the same problems but in different settings (semi-supervised learning and weakly supervised learning) and then discuss what our results mean for textual DA. While Kim et al. (2022); Zhu et al. (2023) findings are relevant to ours, our interest in the results differ. In both papers, the authors attempt to find the best way to use the available data. In our case, we test several settings we think are more realistic, but the point of interest is to find if data augmentation is useful when there is little or no validation data for fine-tuning the classifier. Indeed, there is little doubt that using validation data as training data will be more efficient than the current accepted setting.
6.1.   Evaluating small data learning with more realistic use of data
In academic papers, DA is often evaluated with the assumption that we have access to clean validation data, often in larger quantity than the training data. This validation data is used to fine-tune both the DA algorithm and the classifier, which results in unrealistic settings for practitioners, who may not have access to validation data or want to use their data in better ways. In this section, we evaluate the use of DA in various low-data settings in which we redefine the training and validation split, using different strategies.
We design experiments using two setups, electing to test the setups with the starting size of 10. This is motivated by the long time it takes to fine-tune and run all algorithms [^8] Depending on the final training set size and the DA algorithms, running one experiment takes between 30 and 75 minutes, excluding the grid-search for the hyperparameters, on a RTX 3090., and the fact that the efficiency of DA was the same for both starting sizes, making duplication of the experiments unnecessary. An open question is how these settings would work if we had more initial training data, but we leave this for future work.
Firstly, we assume the validation data is not available (noVal), and we only have training data. In this setting, we train for a random number of epochs between 50 and 150, which is the range for which our models performed best in small data learning.
Secondly, we assume the presence of validation data but redefine the training/validation split, by either 1- using it all as training data and training for a random number of epochs between 4 and 8 (valAsTrain), or 2- keeping some of it as validation data (10, 250) to fine-tune both DA algorithms and the classifier (10 val/250 val).
Overall, this gives us four settings to test, illustrated in Figure 3. We select the following algorithms : Copy, Perfect, EDA, AEDA, BT, CBERT, BART, T5, and GPT3.5-ZS. We use a limited selection for efficiency and because the goal of these experiments is simply to establish whether DA helps at all, rather than which one works best.
Finally, to keep in line with other data augmentation studies, we use a ratio of 10 when the final training size is small (noVal) and of 1 for the other three settings.
@@ 6.1.   Evaluating small data learning with more realistic use of data
Results are presented in table 3, and several important observations can be made. First, in almost all cases and contrary to the main results of this paper, data augmentation reveals itself to be useful. However, the fact that the “Copy” strategy often outperforms the other reinforces our belief that the use of data augmentation is simply to facilitate fine-tuning the network.
@@
Second, there is no significant difference in performance between having 10 and 250 data points as validation data, with the rest as training set. This could be related to diminished increase in model performance as dataset size augment, or point that there is a trade-off in validation/training size and that a validation set of 10 data points is too small to be of use.
@@
Third, we confirm here that methods adding something akin to external data (GPT3.5-ZS and GPT3.5-2S) are the most useful techniques, but only on small training sizes. On larger training sizes, they don’t bring any improvement whatsoever to the model, most likely due to the general inefficiency of data augmentation in this setting. [^9] This is concordant with the literature and our own experiments. We also note that while the performance of GPT3.5-ZS is higher than GPT3.5-3S for two settings, this is due exclusively to its poor performance on the Irony datasets, which we discuss in Section 6.2. For the other three datasets, GPT3.5-3S outperforms GPT3.5-ZS by a large margin.
@@
Finally, it seems that the best strategy overall when the total amount of available data is only a few hundred is to use all of them as training data, with the Copy strategy. This outperforms by a large margin the strategies where we keep some data as validation data, most likely because these strategies overfit. As mentioned, however, there is a large gap between 10 and 250 data and it is possible that some amount of data in the middle would perform best. We leave this for future work.
6.2.   Analysis of the VLLMs
@@
Our main results showed that in a small data setting, 3-shot generation was often more efficient than zero-shot generation, by as much as 18% in some cases. However, this does not hold for the two irony datasets, for which zero-shot outperforms 3-shot. In this section, we analyze the output of the algorithms, in the hope of understanding better why some prompt mechanisms work well on some datasets but not on others.
@@
We first show examples of generated sentences in Tables 4 and 5. For the sake of space, we only show generated sentences for the classes of SST-2 and Irony, but we give all sentences generated by ChatGPT in the associated Github.
Negative
@@
Dataset apparently reassembled from the cutting-room floor of any given daytime soap.
@@
Positive
Dataset béart and berling are both superb , while huppert … is magnificent.
@@
Non-Ironic
Dataset Couldn’t have asked for a better night.
GPT3.5-P Could not have hoped for a more perfect evening.
GPT3.5-ZS Wow, I absolutely love it when my phone battery dies in the middle of an important call. Just what I needed!
GPT3.5-3S Enjoying a relaxing day at the beach with clear skies and warm sunshine. #PerfectDay
Ironic
Dataset work should be fun today :unamused_face:
GPT3.5 - P Today at work is expected to be unenjoyable :unamused_face:
GPT3.5-ZS Oh, how wonderful! Another day of unexpected surprises and setbacks. It’s like the universe’s way of keeping me on my toes!
GPT3.5-3S Declared I’d be more organized and lost my keys for the umpteenth time. I’m the queen of tidiness! :key::mag: #OrganizationQueen
Table 5: Examples of generated sentences for the Irony dataset.
By observing generated sentences, we see that ChatGPT has difficulty understanding the concept of Irony, as well as sticking to the more familiar language of Twitter. As such, adding more sentences doesn’t help because the new sentences are too far from the training distribution to bring valuable information. Nevertheless, the 3-shot strategy does seem to bring the generated sentences closer to the training distribution, as we can at least observe that the new sentences contain hashtags and emojis.
7.   Discussion
In this paper, we test data augmentation for sentence classification and show that in both medium and small data learning, performances of DA had been overestimated by the use of inadequate training protocols. Furthermore, while we have shown that the classical methods of DA are inefficient, we have also demonstrated that simulating external data collection with VLLMs does improve the performance, and future work should therefore focus on this rather than techniques modifying the sentences.
We want to emphasize that our results are only looking at balanced classification task, and furthermore only at short text classification using BERT as a classifier. While this seems a small field of study, it is one of the most popular in textual DA literature, making our findings significant.
Future work should also evaluate whether the protocols are adapted for fields close to interpretable DA. Notably non-interpretable DA, or DA for unbalanced data, might still be an efficient mean of increasing performances. Other textual tasks might also still benefit from DA, due to different goals. As an example, data augmentation for the task of keyphrase generation (Ray Chowdhury et al., 2022; Garg et al., 2023) aims to encourage the network to generate more keyphrases that are absent from the input. It is more than plausible that data augmentation is useful for this field. As another example, data augmentation is sometime used to improve explainability (Chen and Ji, 2020; Ansari et al., 2023), and protocol should therefore be checked thoroughly in independent fields to ensure that it works as intended.
8.   Conclusion
In this paper, we look at the protocol surrounding data augmentation, and demonstrate that data augmentation is mainly used for easier fine-tuning, an effect that can be replicated by simply giving more care to the fine-tuning of the network, or by using a simple strategy where we duplicate the data. We also consider other, more realistic applications of DA, where validation data is either not available or used as training data, instead of keeping hundreds of examples for validation and only a dozen for training the model.
These findings are important as they answer many questions in textual DA which have been standing for years, namely which data augmentation algorithms is best (they all perform similarly except for data generation), why does it help (gives more time to the network to learn), and what makes a generated sentence informative (they don’t bring contradictory information to the network). As noted, data generation remains one of the best way to perform data augmentation, if one can get the VLLMs to generate data adequate for the training distribution.
Future work on textual DA should explore ways to bridge this gap and generate data which is more diverse as well as correspond more closely to the real task. Another pertinent exploration that remains to be done is to thoroughly study the difference between Llama2 and the GPT family. In this paper, we used Llama2-13B-Chat and concluded that it was less efficient than the use of ChatGPT, but it remains to be seen if this is true for all sizes of the model. We also used default parameters of ChatGPT, to stay in line with Piedboeuf and Langlais (2023). It is very probable however that other configurations are more efficient, and this should be studied as well. Finally, we attempted to evaluate DA in more realistic settings, by redefining training/validation split, but the testing should be taken further. For example, it is very likely that a jackknife approach would be a better evaluation method for DA in small data setting, and this should be considered in future research on data augmentation.
9.   Limitations
As stated on the paper, this study as a limited, although important, scope of study. We didn’t look at the impact of data augmentation on other languages than English, or other textual task, and while we performed fine-tuning to the best of our knowledge, it is possible that some combinations of hyperparameters that were not explored change the conclusions here. We also omitted an ablation study due to time limitation, but this would provide more insight into the mechanism of training BERT. Other limitations about the models are the limited hyperparameters explored for ChatGPT, the limited models used for Llama2, and the fact that we did not check the results with other classifiers than BERT.
10.   Bibliographical References
[…]
Appendix A Prompts

### v2

@@ 6.1 On the need to better fine-tuning
We first show the inefficiency of the fine-tuning protocol used in past DA studies by reporting the results they obtain without DA and comparing it to our results (without DA as well). We attempt to replicate the starting sizes, changing only the fine-tuning protocol to see the difference a longer patience, label smoothing, and better fine-tuning does. Table 2 presents the results of our experiments compared to those from the literature. Since the use of label smoothing depends on the dataset, we integrate its use as a hyperparameter.
@@
Kumar+Ubani TREC6 79.4 91.6
@@
SST2 Irony IronyB Trec6 SNIPS Average
Perfect 82.3/85.8 62.4/58.7 19.3/33.7 72.5/80.6 94.0/94.9 66.1/70.7
Copy 61.0/68.1 54.4/57.1 25.5/30.3 36.4/49.3 77.7/83.4 51.0/57.6
EDA 63.4/69.0 54.9/58.9 23.8/30.1 35.8/48.0 76.6/83.2 50.9/57.8
AEDA 62.4/67.1 55.4/56.9 23.4/27.4 33.2/45.7 78.0/84.0 50.5/56.2
BT 60.2/66.8 55.1/58.0 24.7/27.9 29.1/51.5 75.0/82.6 48.8/57.4
CBERT 62.5/65.4 56.3/58.5 23.2/28.3 36.4/44.5 76.3/83.4 50.9/56.0
CBART 62.5/65.6 54.5/57.7 23.6/29.0 36.9/47.2 77.0/84.5 50.9/56.8
T5 64.1/67.4 56.3/58.3 24.1/28.1 34.3/48.5 76.7/84.1 51.1/57.2
GPT3.5-P 65.0/68.3 56.6/57.5 22.7/28.3 34.0/40.5 80.2/84.7 51.7/55.9
GPT3.5-ZS 82.0/78.7 56.4/56.1 23.5/24.6 40.0/48.1 84.8/88.7 57.3/59.2
GPT3.5-3S 87.7/76.1 49.2/53.6 20.9/24.6 58.4/63.9 87.2/89.6 60.7/61.5
Llama2-P 66.3/68.6 55.3/57.1 24.6/24.6 33.4/41.6 78.1/82.8 51.6/55.0
Llama2-ZS 74.8/74.0 54.3/57.2 22.2/24.1 37.3/47.3 78.1/84.1 53.4/57.3
Llama2-3S 63.3/64.4 54.1/54.8 21.0/27.2 48.6/55.8 83.3/88.2 54.0/58.1
Table 3: Average metric over 15 runs for the training set sizes of 10 (left) and 20 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Results for which the difference with the baseline was found to be statistically significant according to a t-test are underlined.
@@
As we can see, carefully fine-tuning the classifier increases its performance significantly, gaining between 2 and 10%. Those gains are actually larger than the ones reported by authors while deploying DA. This aligns well with our hypothesis that the gains observed in previous studies stem only from giving more time for the model to learn. Further, if the DA strategies involve simple transformations of the sentences, it is likely that BERT is not learning new information from these transformations, and that the same result will be achieved with the Copy strategy, which we analyze in the next section.
6.2 On the inefficiency of classical DA
@@
We now turn to DA algorithms, showing that better fine-tuning leads to essentially useless (classical) DA algorithms. Indeed we argue that alleged gains from DA that were shown in the literature can be explained by inadequate training of the (initial) classifier. Results are presented in Table 3.
@@
For the small data setting, we notice that, on average, only two strategies really outperform the baseline : zero-shot and 3-shot generation (both using LLMs), 3-shot generation outperforming zero-shot, but not on all datasets. We explore this phenomenon in Section 6.4. The results for medium sizes (500 and 1000) are presented in Appendix C, and are concordant with more recent literature showing a lack of improvement on such a setting.
An important thing to note is that while some algorithms perform statistically better than the baseline at times, so does the “Copy” strategy. As the duplication of data should bring no new information, this suggests that the training protocol in previous studies was suboptimal and that the gains observed are not due to data augmentation.
@@
One avenue that could be explored to bridge the gap would be to use sampling with replacement while fine-tuning BERT, which should be equivalent to duplication of data. [^9] There would still be minor differences as sampling with replacement would not represent each sample equally, but this difference should reasonably not have a large impact.
@@
Figure 3 provides the results of the statistical tests we conducted. To simplify reading, we group together all tests between the same algorithms, meaning all tests between algorithm A and algorithm B are compiled together. Each entry in the heatmap is therefore composed of 10 t-tests (5 datasets * 2 starting sizes), and we report the percentage of t-tests for which distributions were found to be statistically different, counting only the entries for which the row algorithm performs better than the column algorithm. This gives, for each row, the measure of how many times it beats the algorithm in the column by a statistically significant margin. [^10] We provide the code along with the full saved results for all experiments, as well as the code necessary to run all statistical tests. See Figure 4 in Appendix for medium data settings (500/1000).
Baseline Perfect Copy EDA AEDA BT CBERT CBART T5 GPT3.5-P GPT3.5-ZS GPT3.5-3S Llama2-P Llama2-ZS Llama2-3S
Baseline 0 20 10 10 10 30 10 10 10 20 20 40 10 30 30
Perfect 70 0 70 70 70 70 70 70 70 70 70 80 80 80 80
Copy 20 10 0 0 10 20 10 0 0 30 10 40 20 20 30
EDA 20 10 20 0 10 20 0 0 0 10 20 40 20 10 20
AEDA 10 10 0 0 0 20 0 0 0 10 0 30 20 0 0
BT 20 10 0 0 10 0 10 10 0 10 10 30 10 10 20
CBERT 20 10 10 0 10 10 0 10 0 10 0 20 10 0 20
CBART 10 10 0 0 0 10 0 0 0 10 0 30 10 10 10
T5 20 10 0 0 0 20 10 0 0 10 0 40 10 10 20
GPT3.5-P 30 10 10 20 10 40 20 20 10 0 10 30 10 20 20
GPT3.5-ZS 70 10 60 50 50 50 60 50 50 60 0 40 60 50 30
GPT3.5-3S 60 10 60 60 60 60 60 60 60 60 40 0 60 60 50
Llama2-P 20 10 10 0 10 20 20 10 0 0 0 30 0 0 10
Llama2-ZS 30 10 20 20 20 30 20 20 20 30 0 20 40 0 30
Llama2-3S 40 0 40 40 40 40 40 40 40 40 20 20 40 40 0
@@
From the statistical tests, we note that only three methods outperform the others by a significant and consistent margin : the perfect strategy, in which we simulate collecting and annotating external data, and the zero/k-shot generation with LLMs. All others perform equivalently and do not beat the baseline most of the time, or at least not by a significant margin. This suggests that the slight random variations we observe, which could be taken as a sign that some algorithms perform better than others, are due to randomness of the network and not a superior performance.
6.3 On more realistic uses of data
As mentioned, the second point of interest is to evaluate DA in a more realistic setting with regard to the use of validation data. We select the following algorithms : Copy, Perfect, EDA, AEDA, BT, CBERT, BART, T5, GPT3.5-ZS, and GPT3.5-3S. We use this limited selection for efficiency reasons, but also because our goal here is to establish whether DA helps at all, rather than knowing which DA technique works best. Finally, to keep in line with other DA studies, we use a ratio of 10 when the final training size is small (noVal) and of 1 for the other settings. Results are presented in table 4, and several important observations can be made.
@@ 6.3 On more realistic uses of data
First, in almost all cases and contrary to the main results of this paper, data augmentation reveals itself to be useful. However, the fact that the “Copy” strategy often outperforms the other strategies reinforces our belief that the use of data augmentation is simply to facilitate fine-tuning the network.
@@
Second, there is no difference in performance between having 10 and 250 data points as validation data, with the rest as training set. This could be related to diminished increase in model performance as dataset size augments, or point to a trade-off in validation/training size and that a validation set of 10 data points is too small to be of use.
@@
Third, we confirm here that methods adding something akin to external data (GPT3.5-ZS and GPT3.5-3S) are the most useful techniques, but only on small training sizes. On larger ones, those methods do not bring any improvement whatsoever, most likely due to the general inefficiency of data augmentation in this setting. [^11] This is concordant with the literature and our own experiments. We also note that while the performance of GPT3.5-ZS is higher than GPT3.5-3S for two settings, this is due exclusively to its poor performance on the Irony datasets, which we discuss in Section 6.4. For the other datasets, GPT3.5-3S outperforms GPT3.5-ZS by a large margin.
@@
Finally, it seems that the best strategy overall when the total amount of available data is only a few hundred is to use all of them as training data, with the Copy strategy. [^12] Or, we can assume, a larger number of epoch. This outperforms by a large margin the strategies where we keep some data as validation data, most likely because these strategies overfit. As mentioned, however, there is a large gap between 10 and 250 data and it is possible that some amount of data in the middle would perform better. We leave this investigation for future work.
6.4 Analysis of DA with LLMs
@@
Our main results show that in a small data setting, 3-shot generation was often more efficient than zero-shot generation, by as much as 18% in some cases. However, this does not hold for the two irony datasets, for which zero-shot outperforms 3-shot. In this section, we analyze the output of the algorithms, in the hope of understanding better why some prompt mechanisms work well on some datasets but not on others.
@@
We first show examples of generated sentences in Table 5 and in Appendix A. For the sake of space, we only show generated sentences for the tasks of SST-2 and Irony, but we give all sentences generated by ChatGPT in the associated Github.
Couldn’t have asked for a better night.
-P Could not have hoped for a more perfect evening.
Non-Ironic -ZS Wow, I absolutely love it when my phone battery dies in the middle of an important call. Just what I needed!
-3S Enjoying a relaxing day at the beach with clear skies and warm sunshine. #PerfectDay
Work should be fun today :unamused_face:
-P Today at work is expected to be unenjoyable :unamused_face:
Ironic -ZS Oh, how wonderful! Another day of unexpected surprises and setbacks. It’s like the universe’s way of keeping me on my toes!
-3S Declared I’d be more organized and lost my keys for the umpteenth time. I’m the queen of tidiness! :key::mag: #OrganizationQueen
Table 5: Generated sentences by the GPT3.5-based DA methods for the Irony dataset for one ironic and one non-ironic example.
By observing generated sentences, we see that ChatGPT has difficulty understanding the concept of Irony, as well as sticking to the more familiar language of Twitter. As such, adding more sentences doesn’t help because the new sentences are too far from the training distribution to bring valuable information. Nevertheless, the 3-shot strategy does seem to bring the generated sentences closer to the training distribution, as we can at least observe that the new sentences contain hashtags and emojis. It is likely that by further fine-tuning the prompts, we would reduce the gap between the data distribution and the generated sentences and generate highly informative sentences.
7 Conclusion
In this paper, we test DA for sentence classification and show that in both medium and small data learning, performances of DA had been overestimated by the use of inadequate training protocols. Furthermore, while we have shown that classical DA methods are inefficient, we have also demonstrated that simulating external data collection with LLMs does improve the performance, and future work should therefore focus on this rather than techniques modifying genuine sentences.
We want to emphasize that our results are only looking at balanced classification tasks, and furthermore only at short text classification using BERT as a classifier. While this seems a small field of study, it is one of the most popular in textual DA literature, making our findings significant.
Future work should also evaluate whether the protocols are adapted for fields close to interpretable DA. Notably non-interpretable DA, or DA for unbalanced data, might still be an efficient mean of increasing performances. Other textual tasks might also still benefit from DA, due to different goals. As an example, data augmentation for the task of keyphrase generation (Ray Chowdhury et al., 2022; Garg et al., 2023) aims to encourage the network to generate more keyphrases that are absent from the input. It is very plausible that data augmentation may be useful here.
Our findings are important as they answer many questions in textual DA which have been standing for years, namely which data augmentation algorithms is best (they all perform similarly except for data generation), why does it help (gives more time to the network to learn), and what makes a generated sentence informative (they don’t bring contradictory information to the network). As noted, data generation remains one of the best way to perform data augmentation, if one can get the LLMs to generate data adequate for the training distribution.
8 Limitations
As already stated, this study has a specific, although important, scope. We didn’t look at the impact of data augmentation on other languages than English, or other textual tasks, and while we performed fine-tuning to the best of our knowledge, it is possible that some combinations of hyperparameters that were not explored may change some of our conclusions. We also omitted an ablation study of the fine-tuning (patience, label smoothing, grid search) due to time limitation, but this would provide more insight into the mechanism of training BERT. Furthermore, our exploration of hyperparameters for ChatGPT (mainly the temperature) was limited, and we only used the Llama2-13b-chat model as a comparison, but a comparison of other sizes and also other LLMs should be included in future work, such as Llama3 Meta (2024), Gemini Team et al. (2024), GPT-4 OpenAI et al. (2024), and others. Finally, we only looked at BERT as a classifier, which is the most common one used in DA studies, but it would be pertinent to not only look at other pretrained transformers but also at the impact of DA on more classical models such as CNNs, RNNs, or even statistical methods (SVM, NB, etc).
On a broader scope, there is no guarantee our results hold for other textual tasks (such as question answering, explainability, or keyphrase generation), which have different structures. A final point is that we only compared ourselves to the studies of Kumar et al. (2021); Piedboeuf and Langlais (2023) to demonstrate the inefficiency of fine-tuning in past studies. While these are some of the most complete we found, it would be pertinent to repeat the experience with other papers.
Appendix A Generated sentences
@@
Negative Dataset apparently reassembled from the cutting-room floor of any given daytime soap.
@@
Positive Dataset béart and berling are both superb , while huppert … is magnificent.
@@
Appendix B Prompts